Sheldon owns a mango farm and the weight of his mangoes are

normally distributed with an average weight of 200 grams and a standard

deviation of 15 grams. If there are 800 mangoes in the farm, what is the mean

and standard deviation of the sample mean weights of 50 mangoes?


1
Expert's answer
2021-04-28T05:55:23-0400

Mean: "\bar x=200\;g."

Standard deviation: "s=\frac{15}{\sqrt{50}}\sqrt{\frac{800-50}{800-1}}=2.0552."